Can you help find the missing number?
 
6, 20, 36, 48, 50, ?, 0
 
The answer apparently is 36 (NTSE 2013, First Level MAT), but I’m not sure how it was derived.

Grade:

1 Answers

Ashwin
9 years ago
HI,
 
6,20,36,48,50,?,0
you can write 6 as 1 * 1*6 
20 = 2*2*5
36 = 3*3* 4
48 = 4*4* 3
50 = 5*5*2
? = 6*6*1
0 = 7*7*0
 
therefore ? = 6*6 = 36 
hence 36 is the answer